News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 4 January 2002 9.00am EST

Happy New Year from the Chandra Team!

During the last two weeks the observing schedule was interrupted once following a high radiation event at 1:30am EST on Dec 26 that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and radiation levels dropped sufficiently to allow the loads to be restarted at 6pm on Dec 27. Observations of Beta Ceti, G78.2+2.1, CYG X-1, A168 were impacted by the event and will be rescheduled in a later load.

Preparations are complete for the Jan 6-10 American Astronomical Society (AAS) meeting in Washington DC, where two Chandra press releases are planned.

The schedule of observations for the coming week are shown below and include an observation of XTE J1118+480 coordinated with HST.

------------------------------------------
Radiation Belts Jan 7 M31-T4 ACIS-I NGC507 ACIS-I Jan 8 NGC1550-A ACIS-I NGC1550-B ACIS-I MG J0414+0534 ACIS-S Radiation Belts Jan 9 NGC5866 ACIS-S Jan 10 PSS1326+0743 ACIS-S HS 1111+4033 ACIS-S XTE J1118+480 ACIS-S Jan 10 24 Ursae Majoris ACIS-S/HETG Jan 11 Radiation Belts Jan 12 GX 301-2 ACIS-S/HETG Jan 13 PSR B0833-45 HRC-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports